OpenStudy (anonymous):

evaluate each polynomial for x=3 and x=-3 \[4x ^{2}-6x+9\]

OpenStudy (anonymous):

Simply plug in x = 3, and x = -3 into the equations, and then do the math.

OpenStudy (anonymous):

i did am not sure if i did the whole thing right

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

4*(3)^2 - 6*3 + 9 =

OpenStudy (anonymous):

but how did you get three i thought it was 4(2)^2-6(2)+9

OpenStudy (anonymous):

i see i put 2 instead of 3

OpenStudy (anonymous):

you're letting x=3 so wherever there's an x, replace it with a 3...

OpenStudy (anonymous):

so would it b 12^2-12+9

OpenStudy (anonymous):

no, do the exponentiation first 3^2... then multiply by 4

OpenStudy (anonymous):

awww cute baby...

OpenStudy (anonymous):

thank you and sorry but am still not getting it am a little off when it come to math

OpenStudy (anonymous):

4x^2-6x+9 (a) x=3 4*(3)^2-6*3+9= 4*9-18+9=36-9=27 (b) x=-3 4*(-3)^2-6*(-3)+9=4*9+18+18=36+36=72 Putting x=3 and x=-3 is not giving 0 , hence they are not the roots of this equation.

OpenStudy (anonymous):

there's this thing called order of operations... exponentiation needs to be done before multiplication. so 3^2 is done first: 3^2 = 9 so your now doing this: 4*9 - 6*3 + 9

OpenStudy (anonymous):

ok, you got the answers..

OpenStudy (anonymous):

are you talking to me ? @ dpalnc

OpenStudy (anonymous):

nope, the asker...

OpenStudy (anonymous):

thank you guys had a couple question i didnt understand.

OpenStudy (anonymous):

ooopss , sorry ! :p